Find empirical formula for a sugar with 44.77%C, 7.52%H, 47.71%O.
1
Expert's answer
2020-12-09T06:32:56-0500

Let us assume that 100 g of sugar contain 44.77 g C, 7.52 g H and 47.71 g O. Their amounts in moles should be determined next:


"n(C)=\frac{44.77g}{12.01g/mol}=3.728 mol"


"n(H)=\frac{7.52g}{1.01g/mol}=7.446mol"


"n(O)=\frac{47.71g}{16.00g/mol}=2.982 mol"

Dividing all these values by the lowest of them (2.982) gives the C:H:O ratio of 1.25 : 2.5 : 1. The lowest whole number ratio will be 5:10:4, and the empirical formula is C5H10O4.

Answer: C5H10O4